OpenStudy (anonymous):

What is the range of the function f(x) = 3.2x for the domain {-4, -2, 0, 2, 4}?

OpenStudy (anonymous):

Solve f(x) = 3.2x for each value in our domain f(-4) = 3.2(-4) = f(-2) = 3.2(-2) = f(0) = 3.2(0) = f(2) = 3.2(2) = f(4) = 3.2(4) =

OpenStudy (anonymous):

so i just multiply 3.2 for each of the equations

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Mention me (@mathway) if you need assistance. :)

OpenStudy (anonymous):

ok i got that one but i need help with a second What is the range of the function f(x) = 12 − 3x for the domain {-4, -2, 0, 2, 4}?

OpenStudy (anonymous):

Do the same. Just exchange plug in the domain in x. Like, f(-4)=12-3(-4) and so on.

OpenStudy (anonymous):

Would really appreciate it if you would be a fan and give me a medal. :)